OpenStudy (anonymous):

Use the result that ∫ ℮-a(2)x(2)dx=√∏∕a and √∏=√⅟2 And change of variables techniques to derive the chi-square distribution

OpenStudy (unklerhaukus):

Have i got this equation right?\[\int\limits_0^x e^{-a^2x^2}\text dx=\sqrt{\frac\pi a}\] the second one cant be right\[\sqrt\pi\neq\sqrt\frac 12\]

OpenStudy (anonymous):

Yes Unkle that's the equation.

OpenStudy (unklerhaukus):

how can the square root of pi be the square root of 1/2 pi is not 1/2? am i missing something?

OpenStudy (anonymous):

please Unkle do you see anyway to handle it?

OpenStudy (unklerhaukus):

i was kinda hoping that was typo, the chi-square distribution is the same as the normal or gaussian distribution right?

OpenStudy (unklerhaukus):